OpenStudy (anonymous):

Evaluate this expression. 15 (-3) + 5| -2 | (-5) =

OpenStudy (anonymous):

15 * -3 = -45 abs(-2) = 2 -45 + 10(-5) = -45 + -50 = -95

OpenStudy (anonymous):

(-3) + 5| -2 | (-5) = (-3) + 10 (-5) = 7 X -5= -35

OpenStudy (anonymous):

Wait is 15 the problem or part of the problem?

OpenStudy (anonymous):

part of the problem

OpenStudy (anonymous):

@countonme123 good call

OpenStudy (anonymous):

15/-35 =-3/7